OpenStudy (anonymous):

Find the slope and the y-intercept of the line. 5x=4y+12 The slope is ?

OpenStudy (saifoo.khan):

bring this equation to this form: y = mx + b

OpenStudy (anonymous):

convert it into y=mx+c, where m is the slope of the line and c is the y intercept

OpenStudy (anonymous):

-4y=-5x+12 y=(5/4)x-3 The slope is: (5/4)

OpenStudy (anonymous):

The y-intercept is (0,?)

OpenStudy (saifoo.khan):

Compare: y=m x + c y=(5/4)x-3

OpenStudy (anonymous):

OMG, I am so dumb

OpenStudy (saifoo.khan):

Lol.

OpenStudy (anonymous):

so the y intercept is 4(x) -3?

OpenStudy (anonymous):

No, that can't be right

OpenStudy (anonymous):

(0,-3)

OpenStudy (saifoo.khan):

congrats, (0, -3) is 100% correct.

OpenStudy (anonymous):

NO WAY????

OpenStudy (saifoo.khan):

lol, srsly thats the answer.

OpenStudy (anonymous):

Damn I am good. NOT...LOL

OpenStudy (saifoo.khan):

Lol, now im Jealous, jkjk.